OpenStudy (anonymous):

From midnight to 6:00 am, the temputure rose 8c. at 6:00sm, the temputure was -20c. what was the temputure at midnight. (will give medal) @kx2bay

OpenStudy (anonymous):

-28

OpenStudy (anonymous):

why? -28 + 8 = 20

OpenStudy (anonymous):

wait so what did was -20 - -8 equals -28 right

OpenStudy (anonymous):

no I did x + 8 = -20 x = -20-8

OpenStudy (anonymous):

x been the temp at midnight

OpenStudy (anonymous):

ok so the answer is -28 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it was colder at midnight than at 6am by 8 degrees